OpenStudy (anonymous):

Please verify : assume x,y and z are positive use the properties of logarithms to write -2logb(x)-6logb(y)+1/7logb(z). As one logarithm Logb(x)^-2+ logb(y)^-6 + logb(z)^1/7 Logx^-2y^-6z^1/7?

OpenStudy (anonymous):

\[-2\log_{b} x-6\log_{b} y+\frac{ 1 }{ 7}\log_{b} z=-\log_{b} x^2-\log_{b} y^6+\log_{b} z ^{\frac{ 1 }{ 7 }}\] \[=\log_{b} \frac{ z ^{\frac{ 1 }{ 7 }} }{ x^2y^6 }\]

OpenStudy (anonymous):

I see where I went wrong. Thank you for the help!

OpenStudy (anonymous):

yw